OpenStudy (anonymous):

solve the following equation by using square roots, (x+10)^2=81

OpenStudy (anonymous):

Take the squareroot of both sides. The left side will have x + 10. the right side will be 9 then you will have the equation x + 10 = 9. I think you can solve it from here

OpenStudy (zarkon):

\[x+10=\pm 9\]

OpenStudy (anonymous):

thank you so the answer is x+10=+/- 9

OpenStudy (zarkon):

\[x=-10\pm 9\] x=-1 or -19